Find the vector equation of the plane determined by the points (1, 0, −1), (0, 1, 1) and (−1, 1, 0). Also nd the point of intersection of the line
1
Expert's answer
2020-03-06T16:02:14-0500

Solution:

The equation of a plane passing through three points

"\begin{vmatrix}
 x-1 & y-0&z+1 \\
 0-1 & 1-0&1-(-1)\\
-1-1 & 1-0&1+0
\end{vmatrix}=0,
\\
\begin{vmatrix}
 x-1 & y&z+1 \\
 -1 & 1&2\\
-2 & 1&1
\end{vmatrix}=0,
\\
x-1-z-1-4y+2z+2+y-2x+2=0,"

"-x-3y+z+2=0" ,

"x+3y-z-2=0" .

Answer:

"x+3y-z-2=0" .